OpenStudy (jball96):

Given the equation, -5|x + 1| = -10, the isolated absolute value equation would be |x + 1| = -15.

OpenStudy (mitu12):

i believe it would be Ix+1I=-5

OpenStudy (mitu12):

-10 +5=-5

OpenStudy (skullpatrol):

$$\huge -5|x + 1| = -10$$ Dividing both sides by -5 will isolate the absolute value: $$\huge \dfrac{-5|x + 1|}{-5}= \dfrac{-10}{-5}$$ $$\huge \dfrac{ \cancel{-5}|x + 1|}{\cancel{-5}}= 2$$ $$\huge |x + 1| = 2$$

OpenStudy (agent0smith):

@mitu12 the -5 in \(\Large -5|x + 1| = -10\) is being multiplied by the \(\Large |x + 1|\), so you cannot cancel it out by adding 5; addition can not undo multiplication, only division can.

OpenStudy (skullpatrol):

With the exception that multiplication by 0 cannot be "undone" by division by 0 because division by 0 is undefined.